# Robust Predictive Motion Planning by Learning Obstacle Uncertainty

### v1

@@ IV-D Performance Evaluation
This subsection evaluates the proposed method for learning the intended control set of a system, and the accuracy of the resultant forward reachability analysis.
@@ IV-D1 Comparison Between Different Methods for Learning the Intended Control Set
The performance of formulation (13) for learning the intended control set of the obstacle is compared with the original formulation (8), the moving-horizon (MH) method that fixes $|\mathcal{I}_{t}^{s}|=L$ (see Section IV-B1), and the rigid tube (RT) method [21]. Suppose that the intended control set of a dynamic system is learned by observing the longitudinal and lateral accelerations in the ground coordinate system. The system takes the control action from a bounded but unknown set $\tilde{\mathbb{U}}_{t}^{s}$ with unknown distribution, and its admissible control set $\mathbb{U}_{t}^{s}$ is overly estimated as a regular hexagon. The results are shown in Fig. 3, and the computation time is summarized in Table I. The analyses through Fig. 3 and Table I reflect that the result by (13) closely aligns with the original formulation (8), yet requires less computation time. Compared to the RT method that tends to overestimate certain areas, the learned set by (13) is less conservative. It also surpasses the accuracy of the MH method, which tends to underestimate the set. Notably, the learned set expands when the number of samples increases, while the computation time remains in the same order. This implies an enhanced robustness of the result without compromising computational performance.
IV-D2 Analysis of the Responsiveness of the Method
@@
This subsection analyzes the responsiveness of (13) in learning the control set of a system as in Section IV-D1. The system is designed to perform mild actions before time step $t=50$, and from time step $t=51$ it shifts to executing more aggressive maneuvers from an unknown set $\tilde{\mathbb{U}}_{t}^{s}$. The volume of the learned set with the discrete time steps, as well as the 2-norm of the sampled control action ($||u_{t}^{s}||$), are presented in Fig. 4. In addition, the learned control sets at different time steps are shown in Fig. 5. It is seen in Fig. 4 that after the system starts to take aggressive control actions from $t=51$, and the learned control set can rapidly respond to the change, the delay between the set volume and the change of the control style is just $3$ steps. Fig. 5 shows how the learned set starts from a minor initial set and grows when more samples are obtained. The set expands notably from $t=50$ to $t=54$ as a few aggressive control actions are sampled. The set at $t=100$ contains both the mild and aggressive control actions. This reflects that the method does not lose any information during the online learning process.
Figure 4: Control actions of the system and volume of learned control set.
@@
Figure 5: Learned control sets at different time steps (at $t=50$ the mild control behavior ends, and at $t=100$ the sampling ends).
Figure 6: Predicted occupancy of a nonlinear vehicle model.
@@ IV-D3 Reachability Analysis
This section evaluates the results of reachability analysis through in (IV-C) based on the learned control set $\hat{\mathbb{U}}_{t}^{s}$. The system in this case is modeled by a nonlinear kinematic single-track vehicle model as in (III.1), where the control inputs generated from an unknown input set $\tilde{\mathbb{U}}_{t}^{s}$ with unknown distribution. The accelerations in the global coordinate system are extracted through historical trajectories to learn its intended control set by solving (13). Then it uses (IV-C) to predict the occupancy over a horizon, where the dynamics of the system are modeled by a second-order integrator as in (4). It is seen in Fig. 6 that only a few samples are outside of the predicted occupancy, and the occupancy successfully covers the sampled trajectories at the end of the horizon. This indicates enhanced robustness over time. In addition, even though the nonlinear model of the system and the linear second-order integrator model used for predicting the occupancy in (IV-C) are quite different, the LP-based method still offers the desired accuracy in predicting the occupancy.
@@
In addition, compared with a similar method, our proposed method has advantages in calculation time. Among the related work in Section II-A, the reachability analysis based on solving mixed-integer linear programming (MILP) problem [18] exhibits the most similar characteristics to the proposed method as it does not require the distribution of behaviors and model information of the system. The MILP-based method can get similar occupancy as in Fig. 6 by tuning the confidence level [18, Fig. 4]. For the same prediction horizon, the computation time by the proposed method with $100$ trajectories and $500$ trajectories is $0.0296\ {\rm s}$ and $0.0304\ {\rm s}$. For the MILP-based method, as presented in [18, Table I], the average computation time with $100$ trajectories and $500$ trajectories is $0.0322\ {\rm s}$ and $0.1352\ {\rm s}$. This reflects that our proposed method has a better computational performance than the MILP-based method, and the computation time remains unaffected by the number of samples.
@@ V Robust Motion-Planning Strategy
where $i=1,\ldots,N$. In the cost function (15a), ${r}$ denotes the reference state. The cost function ${V}(\cdot)$ is defined for measuring the performance of the motion planning. The constraints (15b)–(15d) concerning the ego system have been introduced in (1). The constraint (15e) refers to the safety requirement on the distance between the ego system and obstacle $s$, where ${\rm dis}(\cdot)$ is a distance measure between $p^{e}_{i|t}$ the position of center of geometry of the ego system and the obstacle’s occupancy $\hat{\mathcal{O}}_{i|t}^{s}$ [34]. It means at time step $t+i$ in the prediction horizon, the distance between $p^{e}_{i|t}$ and $\hat{\mathcal{O}}_{i|t}^{s}$ should be larger than the minimum safety distance $d^{s}_{\rm min}$. Note that $p_{i|t}^{e}$ and $\hat{\mathcal{O}}_{i|t}^{s}$ predict the motion of the center of geometry of the ego and obstacle systems, such that the actual shape factors of them should be considered in designing $d^{s}_{\rm min}$. In addition, (15e) can be designed with a slack variable to enhance the feasibility. This will be introduced in Example V.1 with the explicit and equivalent expressions of constraint (15e).
The motion planning of the ego system is performed by solving the OCP (V) at every time step $t$ to obtain the optimal control sequence $\{u^{e,\star}_{i-1|t}\}_{i=1}^{N}$, which is applied to steer the model (15b) to generate the reference trajectory for a lower-level controller at the current time step.
@@
MPC formulation of the reach-avoid problem: Consider the models of EV and SV in Example III.1. Based on the occupancy $\hat{\mathcal{O}}_{i|t}^{s}$ of the SV, the EV model (III.1), and formulation (V), the OCP in this case is designed as
@@
$\underline{\mathcal{U}}\leq\left[v_{i|t}^{e}\ a_{i|t}^{e}\ \delta_{i-1|t}^{e}\right]^{\rm T}\leq\overline{\mathcal{U}},$ (16c)
@@
$(H_{i|t}^{s}p_{i|t}^{e}-h_{i|t}^{s})^{\rm T}\lambda_{i}^{s}\geq d^{s}_{\rm min}-\varepsilon_{i}^{s},$ (16e)
$||(H_{i|t}^{s})^{\rm T}\lambda_{i}^{s}||_{2}=1,$ (16f)
@@
where $Q_{1}$,normal-$\cdots$,$Q_{4}$ are weighting matrices, (16b) is the discrete approximation of the continuous model (III.1) using a fourth-order Runge-Kutta method with the sampling interval $T$, $u_{i-1|t}^{e}=[\delta_{i-1|t}^{e}\ \eta_{i-1|t}^{e}]$ is the control input of model (III.1), and $x_{i-1|t}^{e}$ is the state of model (III.1). Parameters $\underline{\mathcal{U}}$ and $\overline{\mathcal{U}}$ mean the lower and upper bounds on the velocity, acceleration, and front tire angle of the EV. Constraint (16d) limits the position of the center of geometry of the EV, $p_{i|t}^{e}=[p_{x,i|t}^{e}\ p_{y,i|t}^{e}]^{\rm T}$, within the driveable area $\mathcal{D}$. Constraints (16e)–(16g) are an equivalent reformulation of the collision-avoidance constraint (15e) with additional decision variable $\lambda_{i}^{s}$ and slack variable $\varepsilon_{i}^{s}$ [34, Section III-B]. The parameter $n_{h}$ is the number of edges of the set $\hat{\mathcal{O}}^{s}_{i|t}$, and the matrix $H^{s}_{i|t}$ and the vector $h^{s}_{i|t}$ are given by
@@
Define $l^{e}$ and $w^{e}$ as the length and width of the EV, $l^{s}$ and $w^{s}$ the length and width of the SV. The safety distance $d^{s}_{\rm min}$ in (16e) is designed as
$d^{s}_{\rm min}=\sqrt{(l^{e}/2)^{2}+(w^{e}/2)^{2}}+\sqrt{(l^{s}/2)^{2}+(w^{s}/2)^{2}},$
@@ VI Simulations of Reach-Avoid Motion Planning
This section presents simulation results for reach-avoid planning as shown in Fig. 1, where the EV and SV are specified as a car-like mobile robot in this case. The implementations can be found in our published code [^2] https://github.com/JianZhou1212/robust-mpc-motion-planning-by-learning-obstacle-uncertainties. The mission of the EV is to follow a predefined reference state while avoiding collision with the SV and borders of the driveable area. It is assumed that the SV holds a higher priority, requiring the EV to actively take measures to avoid collisions. The controller of the SV is simulated by a nonlinear MPC to track its reference target, where the reference target and the model and controller parameters are unknown to the EV. The SV takes control input from an intended control set $\tilde{\mathbb{U}}_{t}^{s}$ to plan a reference trajectory to reach its reference target; the set $\tilde{\mathbb{U}}_{t}^{s}$ and the distribution of control actions of the SV are also unknown to the EV. From the perspective of the EV, the SV is modeled by (4).
In the numerical studies, the computation of the occupancy of the SV firstly requires formulating an admissible control set $\mathbb{U}^{s}$ according to Assumption IV.2. Next, the LP problem (13) is solved to learn the unknown intended control set of SV. This is initialized by an initial information set $\mathcal{I}_{0}^{s}$ that contains some artificial samples to construct a small but non-empty set $\hat{\mathbb{U}}_{0}^{s}$ by solving (8). This implies that the proposed method does not have any prior information on the uncertainties of the SV. Then, for $t\in\mathbb{N}_{+}$ the set $\hat{\mathbb{U}}_{t}^{s}$ is obtained by solving (13) based on $\hat{\mathbb{U}}_{t-1}^{s}$ and $u_{t-1}^{s}$, where $\hat{\mathbb{U}}_{t-1}^{s}$ is obtained at time step $t-1$, and $u_{t-1}^{s}$ contains the measured longitudinal and lateral accelerations of the SV in the ground coordinate system at time step $t-1$. The learned set $\hat{\mathbb{U}}_{t}^{s}$ is substituted into (IV-C) to obtain $\hat{\mathcal{O}}_{i|t}^{s}$, i.e., the predicted occupancy of the SV, for planning of the EV by solving (V.1).
The proposed planner is compared with a robust MPC (RMPC) and a deterministic MPC (DMPC), where the RMPC and DMPC are implemented by replacing $\hat{\mathbb{U}}_{t}^{s}$ in (14b) with $\mathbb{U}^{s}$ and ${\bm{0}}$, respectively. The set addition and polytope computations were implemented by the Python package pytope [35]. The optimization problems involved were solved by CasADi [36] and Ipopt [37] using the linear solver MA57 [38]. The simulations were performed on a standard laptop running Ubuntu 22.04 LTS and Python 3.10.12. Table II provides a summary of simulation parameters.
Figure 8: Comparison of $d^{e,s}$ and $d^{e,\rm ref}$ during the motion. (a) The polytope distance between EV and SV ($d^{e,s}$) during the motion (the minimum EV-SV distance with DMPC is $0.03\ {\rm m}$, and that with the proposed method is $0.14\ {\rm m}$). (b). The Euclidean distance between EV and its reference target ($d^{e,\rm ref}$) during the motion. (c). The cost function value.
Planner & success rate ${\rm min}(d^{e,s})$ $\tau^{e,\rm ref}$ Sum of Cost
Name Collision-free rate Complete rate Mean Min. Mean Max. Mean Max.
Proposed $100\%$ $100\%$ $0.104$ $0.062$ $8.75$ $8.75$ $2000$ $2331$
RMPC $100\%$ $73\%$ $0.711$ $0.199$ $11.26$ $13.75$ $4526$ $6357$
DMPC $76\%$ $100\%$ $0.013$ $5.7{\rm e}^{-4}$ $8.75$ $8.75$ $1966$ $1980$
@@
The methods are first implemented in a single case to get insights into the performance of each planner. The comparison is based on the planned paths, the polytope distance between the EV and SV ($d^{e,s}$), the Euclidean distance between the EV and its reference state ($d^{e,\rm ref}$), which reflects the convergence speed to the reference, and the cost function value. The results are shown in Fig. 7 and Fig. 8, respectively. It is observed in Fig. 7 that the three methods perform differently in the same scenario as a result of different considerations of uncertainties. Further analysis from Fig. 8(a) shows that the minimum distance between EV and SV with DMPC is close to $0\ {\rm m}$, and that with the proposed method is $0.14\ {\rm m}$, which is sufficient in this case considering that the length of the EV is $0.26\ {\rm m}$. In contrast, RMPC can generate the safest reference trajectory as it maintains a larger distance between EV and SV. However, the cost of pursuing robustness is that the EV needs a longer time and a larger cost to converge to the reference state, as shown in Figs. 8(b)–(c), while the DMPC and the proposed method converges faster with a smaller cost.

### v2

@@ IV-D Performance Evaluation
This subsection evaluates different methods for learning the intended control set of an obstacle system, and the accuracy of the resultant forward reachability analysis.
@@ IV-D1 Comparison Between Different Methods for Learning the Intended Control Set
The performance of the online recursion (13) and the moving-horizon (MH) method for learning the intended control set of the obstacle is compared with the original formulation (8), a rigid tube (RT) method [22], and a method that directly learns the set as a convex hull of the samples in the information set $\mathcal{I}_{t}^{s}$ (the convex hull is computed using the Python package pytope [35]). Assume that the intended control set of the obstacle is learned by observing the longitudinal and lateral accelerations in the ground coordinate system. The results for learning the sets with different numbers of samples are shown in Fig. 3, where the admissible control set $\mathbb{U}^{s}$ is overly estimated as a square (4 vertices) and a regular hexagon (6 vertices), respectively. Furthermore, Table I compares the computation time for learning the set and reachability prediction based on the learned set as in (14).
The analyses through Fig. 3 and Table I reflect that formulation (13) closely aligns with the original formulation (8), yet requires less computation time. Compared to the RT method, the learned set by (13) is less conservative. The set obtained by the MH method is smaller than that by (13). Notably, Fig. 3 shows that the set expands when the number of samples increases, while Table I shows that the computation time of (13) and the MH method remains in the same order. The convex hull-based approach can more tightly contain the samples. However, it has a more complex geometry than the learned set obtained by other approaches, such that it renders a more complicated prediction of reachability. In addition, since the geometry of the convex hull is not deterministic, this is not beneficial for constructing the collision-avoidance constraint in the motion planner (see Example V.1). In comparison, other approaches can determine the geometry of the learned set by defining the shape of $\mathbb{U}^{s}$, such that the forward reachability prediction is more efficient. One additional observation from Table I is that the shape of $\mathbb{U}^{s}$ can affect the computational efficiency of the reachability prediction. A more complex $\mathbb{U}^{s}$ has more degrees of freedom, while also increasing the computation time of reachability prediction. This implies a trade-off between accuracy and efficiency should be considered in practical applications when defining the admissible set $\mathbb{U}^{s}$.
IV-D2 Responsiveness Analysis
@@
This subsection analyzes the responsiveness of the online recursion formulation (13) and the MH approach in learning the control set of an obstacle as in Section IV-D1. In the setup, the obstacle performs mild actions before time step $t=50$, and from time step $t=51$ it shifts to executing more aggressive maneuvers. The volume of the learned set with the discrete time steps, as well as the 2-norm of the sampled control action ($||u_{t}^{s}||$), are presented in Fig. 4. In addition, the learned control sets at different time steps are shown in Fig. 5. Note that the index of the time step in Fig. 4 is equal to the number of online samples as the method obtains one sample at every time step.
It is seen in Fig. 4 that after the system starts to take aggressive control actions from $t=51$, the learned control sets by (13) and the MH approach can both rapidly respond to the change, the delay between the set volume and the change of the control style is just $1$ step. Fig. 5 shows how the learned sets by two approaches start from a minor initial set and grows when more samples are obtained. The set expands notably from $t=50$ to $t=54$ as a few aggressive control actions are sampled. At the end of sampling $t=100$, the set obtained by (13) contains both the mild and aggressive control actions, while the set obtained by the MH approach only contains the aggressive control actions. This shows that (13) predicts the motion of the obstacle based on all observed control actions of the obstacle system, while the MH approach captures the time-varying intent of the obstacle by discarding past information.
Figure 4: Control actions of the system and volume of the learned set.
@@
Figure 5: Learned control sets at different time steps (at $t=50$ the mild control behavior ends, and at $t=100$ the sampling ends). (a) With the online recursion formulation (13). (b) With the MH approach.
Figure 6: Predicted occupancy of the linear model (4) by the proposed method based on the trajectory samples of a nonlinear vehicle model. (a) By the online recursion formulation (13). (b) By the moving-horizon (MH) approach.
@@ IV-D3 Reachability Analysis
This section evaluates the results of reachability analysis through (14) based on the learned control set $\hat{\mathbb{U}}_{t}^{s}$. The actual obstacle dynamics are in this case modeled by a nonlinear kinematic single-track vehicle model as in (3), where the control inputs are generated from an unknown input set with an unknown distribution. The accelerations in the global coordinate system are extracted through historical trajectories to learn its intended control set by (13) and the MH approach, respectively. Then it uses (14) to predict the occupancy over a horizon, where the dynamics of the system are modeled by a second-order integrator as in (4). It is seen in Fig. 6 that the predicted occupancy using a linear model can well cover the trajectory samples generated by the nonlinear kinematic model, which indicates that the linear model-based prediction by the proposed method is a good local approximation of the nonlinear vehicle model with the choosen prediction horizon. Since the occupancy set in Fig. 6 can not completely cover the trajectory samples, it implies that the proposed method, in theory, does not guarantee safety in motion planning, but instead recovers approximately safe behavior by relying on past data to predict future obstacle dynamics. However, the empirical results in Section VI indicate that the proposed method still exhibits a high safety performance as a result of implementing the motion planner in a receding horizon manner. Note that the range of the occupancy in Fig. 6 depends on the maneuvers of the observed obstacle. If the control actions of the obstacle are generated from a larger input set, the predicted occupancy set will expand accordingly. To conclude, even though the nonlinear model of the obstacle and the linear second-order integrator model used for predicting the occupancy in (14) are quite different, the proposed method still offers the desired accuracy in predicting the occupancy.
@@
In addition, compared with a similar method, our proposed method has advantages in calculation time. Among the related work in Section II-A, the reachability analysis based on solving mixed-integer linear programming (MILP) problem [19] exhibits the most similar characteristics to the proposed method as it does not require the distribution of behaviors and model information of the obstacle. The MILP-based method can get similar occupancy as in Fig. 6 by tuning the confidence level [19, Fig. 4]. For the same horizon, the average computation time by the proposed method is $0.0340\ {\rm s}$ for formulation (13) and $0.0432\ {\rm s}$ for the MH approach, with $100$ trajectories, and $0.0361\ {\rm s}$ for formulation (13) and $0.0451\ {\rm s}$ for the MH approach, with $500$ trajectories. For the MILP-based method, as presented in [19, Table I], the average computation time with $100$ trajectories and $500$ trajectories is $0.0322\ {\rm s}$ and $0.1352\ {\rm s}$. This reflects that the proposed method is more computationally efficient compared with the MILP-based approach, and the computation time remains unaffected by the number of sampled trajectories.
Remark IV.2.
Assumption IV.1 ensures the existence of a unique solution of the control input $u_{t-1}^{s}$ in (2) to form a non-empty information set $\mathcal{I}_{t}^{s}$. Given a non-empty set $\mathcal{I}_{t}^{s}$, the convex optimization problem for determining the control set $\hat{\mathbb{U}}_{t}^{s}$, either by (8) or (13), is feasible since the information set $\mathcal{I}_{t}^{s}$ is always contained within the admissible set $\mathbb{U}^{s}$. In addition, state-measurement uncertainties inherently contribute to uncertainties of $u_{t-1}^{s}$. The set $\hat{\mathbb{U}}_{t}^{s}$ learned with small uncertainties in $u_{t-1}^{s}$ would still be expected to center around the set learned with the true $u_{t-1}^{s}$ values. When these uncertainties are prominent, improving the accuracy of $\hat{\mathbb{U}}_{t}^{s}$ would depend on a more precise tracking system.
Remark IV.3.
For the double-integrator model (4), if the obstacle is observed with a constant velocity, then the estimated control input $u_{t-1}^{s}$ is always $\bm{0}$. In this case, the information set is a singleton, and the proposed motion-prediction method corresponds to a constant velocity model. The proposed method is designed to capture what has been observed from the obstacle. In particular, if the obstacle only moves with a constant velocity, the method is supposed to reflect that behavior. However, as soon as the obstacle changes the driving behavior, the method can adapt to this change with one time step, as indicated in Figs. 45. The information set transitions from a singleton to a non-singleton when a non-zero control input from the obstacle is estimated, thereby improving the robustness of the prediction over the constant velocity model.
@@ V Robust Motion-Planning Strategy
where $i=1,\ldots,N$. In the cost function (15a), ${r}$ denotes the reference state. The cost function ${V}(\cdot)$ is defined for measuring the performance of the motion planning. The constraints (15b)–(15d) concerning the ego system have been introduced in (1). The constraint (15e) refers to the safety requirement on the distance between the ego system and obstacle $s$, where ${\rm dis}(\cdot)$ is a distance measure between $p^{e}_{i|t}$ the position of center of geometry of the ego system and the obstacle’s occupancy $\hat{\mathcal{O}}_{i|t}^{s}$ [36]. It means at time step $t+i$ in the prediction horizon, the distance between $p^{e}_{i|t}$ and $\hat{\mathcal{O}}_{i|t}^{s}$ should be larger than the minimum safety distance $d^{s}_{\rm min}$. Note that $p_{i|t}^{e}$ and $\hat{\mathcal{O}}_{i|t}^{s}$ predict the motion of the center of geometry of the ego and obstacle systems, such that the actual shape factors of them should be considered in designing $d^{s}_{\rm min}$. In addition, a large occupancy $\hat{\mathcal{O}}_{i|t}^{s}$ may cause infeasibility of the motion-planning problem as a result of the limited driving area. This can be handled by introducing a slack variable to (15e) to enhance the feasibility, and this will be enforced in Example V.1 with the explicit and equivalent expressions of constraint (15e).
The motion planning of the ego system is performed by solving the OCP (15) at every time step $t$ to obtain the optimal control sequence $\{u^{e,\star}_{i-1|t}\}_{i=1}^{N}$, which is applied to steer the model (15b) to generate the reference trajectory for a lower-level controller at the current time step.
@@
MPC formulation of the reach-avoid problem: Consider the models of EV and SV in Example III.1. Based on the occupancy $\hat{\mathcal{O}}_{i|t}^{s}$ of the SV, the EV model (3), and formulation (15), the OCP in this case is designed as
@@
$\underline{\mathcal{U}}\leq\left[v_{i|t}^{e}\ a_{i|t}^{e}\ \delta_{i-1|t}^{e}\right]^{\top}\leq\overline{\mathcal{U}},$ (16c)
@@
$(H_{i|t}^{s}p_{i|t}^{e}-h_{i|t}^{s})^{\top}\lambda_{i}^{s}\geq d^{s}_{\rm min}-\varepsilon_{i}^{s},$ (16e)
$||(H_{i|t}^{s})^{\top}\lambda_{i}^{s}||_{2}=1,$ (16f)
@@
where $Q_{1}$$\cdots$,$Q_{4}$ are weighting matrices, (16b) is the discrete approximation of the continuous model (3) using a fourth-order Runge-Kutta method with the sampling interval $T$, $u_{i-1|t}^{e}=[\delta_{i-1|t}^{e}\ \eta_{i-1|t}^{e}]$ is the control input of model (3), and $x_{i-1|t}^{e}$ is the state of model (3). Parameters $\underline{\mathcal{U}}$ and $\overline{\mathcal{U}}$ mean the lower and upper bounds on the velocity, acceleration, and front tire angle of the EV. Constraint (16d) limits the position of the center of geometry of the EV, $p_{i|t}^{e}=[p_{x,i|t}^{e}\ p_{y,i|t}^{e}]^{\top}$, within the driveable area $\mathcal{D}$. Constraints (16e)–(16g) are an equivalent reformulation of the collision-avoidance constraint (15e) with additional decision variable $\lambda_{i}^{s}$ and slack variable $\varepsilon_{i}^{s}$ [36, Section III-B]. In practical applications, the variable $\varepsilon_{i}^{s}$ should be penalized with a high weight to make sure that the hard collision-avoidance constraint is violated only in extreme cases. The parameter $n_{h}$ is the number of edges of the set $\hat{\mathcal{O}}^{s}_{i|t}$, and the matrix $H^{s}_{i|t}$ and the vector $h^{s}_{i|t}$ are given by
@@
Define $l^{e}$ and $w^{e}$ as the length and width of the EV, $l^{s}$ and $w^{s}$ as the length and width of the SV. The safety distance $d^{s}_{\rm min}$ in (16e) is designed as
$d^{s}_{\rm min}=\sqrt{(l^{e}/2)^{2}+(w^{e}/2)^{2}}+\sqrt{(l^{s}/2)^{2}+(w^{s}/2)^{2}}.$ (17)
@@
The formulation (17) is designed to achieve a trade-off between the geometric accuracy in the description of the SV and the computational efficiency of the method. For higher geometric accuracy, it is necessary to predict the reachability with consideration of the orientation of the SV. However, integrating the orientation into the obstacle occupancy will make the obstacle occupancy significantly more complex. This also implies that more decision variables or constraints are needed to formulate the collision-avoidance constraints in the motion planner (see [36, Section IV] and [37, Section IV.A]). To this end, (17) is applied to approximate the geometry of the SV. Although this approach may introduce possible conservatism, the potential feasibility issues are addressed in practical applications by using soft constraints as in (16e).
@@ VI Simulations of Reach-Avoid Motion Planning
$\underline{\mathcal{U}}$ $[-1.5\ -0.5\ -0.3]^{\top}$ $\overline{\mathcal{U}}$ [1.5 0.5 0.3]superscriptdelimited-[]$[1.5\ 0.5\ 0.3]^{\top}$
[…]
This section presents simulation results for reach-avoid planning as shown in Fig. 1, where the EV and SV are specified as a car-like mobile robot in this case. The implementations can be found in our published code [^2] https://github.com/JianZhou1212/robust-mpc-motion-planning-by-learning-obstacle-uncertainties. The mission of the EV is to follow a predefined reference state while avoiding collision with the SV and borders of the driveable area. It is assumed that the SV holds a higher priority, requiring the EV to actively take measures to avoid collisions. The controller of the SV is simulated by a nonlinear MPC to track its reference target, where the reference target and the model and controller parameters are unknown to the EV. The SV takes control input from an unknown set to plan a reference trajectory to reach its reference target. The SV’s dynamics are not available to the EV, from the perspective of the EV, the SV is modeled by the linear model (4).
In the numerical studies, the computation of the occupancy of the SV firstly requires formulating an admissible control set $\mathbb{U}^{s}$ according to Assumption IV.2. Since the simulation runs for just $55$ time steps, the size of the information set $\mathcal{I}_{t}^{s}$ is not overly big. Therefore, the online recursion (13) is more suitable for learning the unknown intended control set of SV in this case. This is initialized by an initial information set $\mathcal{I}_{0}^{s}$ that contains some artificial samples to construct a small but non-empty set $\hat{\mathbb{U}}_{0}^{s}$ by solving (8). This implies that the proposed method does not have any prior information on the uncertainties of the SV. Then, for $t\in\mathbb{N}_{+}$ the set $\hat{\mathbb{U}}_{t}^{s}$ is obtained by solving (13) based on $\hat{\mathbb{U}}_{t-1}^{s}$ and $u_{t-1}^{s}$, where $\hat{\mathbb{U}}_{t-1}^{s}$ is obtained at time step $t-1$, and $u_{t-1}^{s}$ contains the measured longitudinal and lateral accelerations of the SV in the ground coordinate system at time step $t-1$. The learned set $\hat{\mathbb{U}}_{t}^{s}$ is substituted into (14) to obtain $\hat{\mathcal{O}}_{i|t}^{s}$, i.e., the predicted occupancy of the SV, for planning of the EV by solving (16).
The proposed planner is compared with a robust MPC (RMPC) and a deterministic MPC (DMPC), where the RMPC and DMPC are implemented by replacing $\hat{\mathbb{U}}_{t}^{s}$ in (14b) with $\mathbb{U}^{s}$ and ${\bm{0}}$, respectively. For the DMPC, the obstacle system (4) is predicted with ${\bm{0}}$ input, such that the prediction is equivalent to a constant velocity model. The set addition and polytope computations were implemented by the Python package pytope [35]. The optimization problems involved were solved by CasADi [38] and Ipopt [39] using the linear solver MA57 [40]. The simulations were performed on a standard laptop running Ubuntu 22.04 LTS and Python 3.10.12. Table II provides a summary of simulation parameters.
Figure 8: Comparison of $d^{e,s}$, $d^{e,\rm ref}$, and the cost during the motion. (a) The polytope distance between EV and SV ($d^{e,s}$) during the motion (the minimum EV-SV distance with DMPC is $0.03\ {\rm m}$, and that with the proposed method is $0.14\ {\rm m}$). (b). The Euclidean distance between EV and its reference target ($d^{e,\rm ref}$) during the motion. (c). The cost function value.
Name Collision-free rate Complete rate Mean $d_{\rm min}^{e,s}$ Min. $d_{\rm min}^{e,s}$ Mean $\tau^{e,\rm ref}$ Max. $\tau^{e,\rm ref}$ Mean $J_{\rm sum}$ Max. $J_{\rm sum}$
$N=10$ $N=8$ $N=10$ $N=8$ $N=10$ $N=8$ $N=10$ $N=8$ $N=10$ $N=8$ $N=10$ $N=8$ $N=10$ $N=8$ $N=10$ $N=8$
Proposed $100\%$ $100\%$ $100\%$ $100\%$ $0.105\ {\rm m}$ $0.081\ {\rm m}$ $0.057\ {\rm m}$ $0.039\ {\rm m}$ $8.75\ {\rm s}$ $10.5\ {\rm s}$ $8.75\ {\rm s}$ $10.5\ {\rm s}$ $2001$ $2576$ $2301$ $2837$
RMPC $98.3\%$ $100\%$ $65.1\%$ $80\%$ $0.689\ {\rm m}$ $0.184\ {\rm m}$ $0.132\ {\rm m}$ $0.083\ {\rm m}$ $11.3\ {\rm s}$ $11.8\ {\rm s}$ $13.75\ {\rm s}$ $13.25\ {\rm s}$ $4496$ $3452$ $5961$ $4702$
DMPC $40.3\%$ $30.7\%$ $100\%$ $100\%$ $0.020\ {\rm m}$ $0.019\ {\rm m}$ $0.010\ {\rm m}$ $0.010\ {\rm m}$ $8.75\ {\rm s}$ $10.5\ {\rm s}$ $8.75\ {\rm s}$ $10.5\ {\rm s}$ $1967$ $2553$ $1983$ $2680$
@@
TABLE IV: Comparison of Computational Performance of Three Methods
Methods Learn Set (13) Reach. Prediction (14) Optimization (16) Total
$N=10$ $N=8$ $N=10$ $N=8$ $N=10$ $N=8$ $N=10$ $N=8$
Proposed 1.24 $\pm$ 0.806 $\rm{ms}$ 1.23 $\pm$ 0.802 $\rm{ms}$ 48.7 $\pm$ 2.19 $\rm{ms}$ 28.0 $\pm$ 2.56 $\rm{ms}$ 48.4 $\pm$ 15.6 $\rm{ms}$ 42.5 $\pm$ 14.0 $\rm{ms}$ 98.5 $\pm$ 15.8 $\rm{ms}$ 72.0 $\pm$ 14.5 $\rm{ms}$
DMPC None None 5.87 $\pm$ 1.59 $\rm{ms}$ 4.90 $\pm$ 1.63 $\rm{ms}$ 50.7 $\pm$ 21.4 $\rm{ms}$ 44.9 $\pm$ 15.1 $\rm{ms}$ 56.7 $\pm$ 21.5 $\rm{ms}$ 49.8 $\pm$ 15.2 $\rm{ms}$
RMPC None None 48.8 $\pm$ 2.53 $\rm{ms}$ 27.9 $\pm$ 1.97 $\rm{ms}$ 62.4 $\pm$ 29.1 $\rm{ms}$ 42.9 $\pm$ 22.2 $\rm{ms}$ 111.3 $\pm$ 29.3 $\rm{ms}$ 70.9 $\pm$ 22.3 $\rm{ms}$
@@
The methods are first implemented in a single case to get insights into the performance of each planner. The comparison is based on the planned paths, the polytope distance between the EV and SV ($d^{e,s}$), the Euclidean distance between the EV and its reference state ($d^{e,\rm ref}$), which reflects the convergence speed to the reference, and the cost function value. The results are shown in Fig. 7 and Fig. 8, respectively. It is observed in Fig. 7 that the three methods perform differently in the same scenario as a result of different considerations of uncertainties. Further analysis from Fig. 8(a) shows that the minimum distance between EV and SV with DMPC is close to $0\ {\rm m}$, and that with the proposed method is $0.14\ {\rm m}$, which is sufficient in this case considering that the length of the EV is $0.26\ {\rm m}$. In contrast, RMPC can generate the safest reference trajectory as it maintains a larger distance between EV and SV. However, the cost of pursuing robustness is that the EV needs a longer time and a larger cost to converge to the reference state, as shown in Figs. 8(b)–(c), while the DMPC and the proposed method converge faster with a smaller cost.